Evaluate the following:
2 tan^2 45° + cos^2 30° – sin^2 60

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

2 tan^245° + cos^230°–sin^260

We know that, the values of the trigonometric ratios are:sin 60° =√3/2
cos 30° =√3/2
tan 45° = 1

Substitute the values in the given problem
2 tan^245° + cos^230°–sin^260
= 2(1)^2+(√3/2)^2-(√3/2)^2
2 tan^245° + cos^230°–sin^260
= 2 + 0
2 tan^245° + cos^230°–sin^260
= 2

Thanks